# An In-depth Look at Gemini’s Language Abilities

### 18 Dec 2023 v1

@@ Abstract
The recently released Google Gemini class of models are the first to comprehensively report results that rival the OpenAI GPT series across a wide variety of tasks. In this paper, we do an in-depth exploration of Gemini’s language abilities, making two contributions. First, we provide a third-party, objective comparison of the abilities of the OpenAI GPT and Google Gemini models with reproducible code and fully transparent results. Second, we take a closer look at the results, identifying areas where one of the two model classes excels. We perform this analysis over 10 datasets testing a variety of language abilities, including reasoning, answering knowledge-based questions, solving math problems, translating between languages, generating code, and acting as instruction-following agents. From this analysis, we find that Gemini Pro achieves accuracy that is close but slightly inferior to the corresponding GPT 3.5 Turbo on all tasks that we benchmarked. We further provide explanations for some of this under-performance, including failures in mathematical reasoning with many digits, sensitivity to multiple-choice answer ordering, aggressive content filtering, and others. We also identify areas where Gemini demonstrates comparably high performance, including generation into non-English languages, and handling longer and more complex reasoning chains. Code and data for reproduction can be found at https://github.com/neulab/gemini-benchmark
@@ 1 Introduction
Gemini is the most recent in a series of large language models released by Google DeepMind (Gemini Team, 2023). It is notable in particular because the results reported by the Gemini team are the first to rival the OpenAI GPT model series (Brown et al., 2020) across a wide variety of tasks. Specifically, Gemini’s “Ultra” version reportedly outperforms GPT-4 on a wide variety of tasks, while Gemini’s “Pro” version is reportedly comparable to GPT-3.5 Gemini Team (2023). Despite the potential impact of these results, the exact evaluation details and model predictions have not been released, limiting the ability to reproduce, inspect, and analyze the results and their implications in detail.
@@
A summary of our main results can be found in Table 1. In sum, we found that across all tasks, as of this writing (December 18, 2023), Gemini’s Pro model achieved comparable but slightly inferior accuracy compared to the current version of OpenAI’s GPT 3.5 Turbo. In the following sections, we will detail our experimental methodology (Section 2) and then perform an in-depth description and analysis of the results on each task. Each analysis is accompanied by an online results browser using Zeno (Cabrera et al., 2023), [^2] https://zenoml.com which can be accessed through the [Uncaptioned image] images in this PDF. All results and code for reproduction can be found at https://github.com/neulab/gemini-benchmark.
@@
Knowledge-based QA MMLU (5-shot) 64.12 67.75 80.48 -
MMLU (CoT) 60.63 70.07 78.95 -
Reasoning BIG-Bench-Hard 65.58 71.02 83.90 41.76
Mathematics GSM8K 69.67 74.60 92.95 58.45
SVAMP 79.90 82.30 92.50 73.20
ASDIV 81.53 86.69 91.66 74.95
MAWPS 95.33 99.17 98.50 89.83
Code Generation HumanEval 52.44 65.85 73.17 -
ODEX 38.27 42.60 46.01 -
Machine Translation FLORES (0-shot) 29.59 37.50 46.57 -
FLORES (5-shot) 29.00 38.08 48.60 -
Web Agents WebArena 7.09 8.75 15.16 1.37
Table 1: Main results of our benchmarking. The best model is listed in bold, and the second best model is underlined. Mixtral was only evaluated on a subset of the tasks.
@@ 2.1 Models Tested
is the second largest model in the Gemini Series, next to the largest Gemini Ultra. [^3] Gemini Ultra is not yet publicly available, and thus we do not test it in the current version of this paper. The model is based on the Transformer (Vaswani et al., 2017) architecture and was trained multimodaly over videos, text, and images. The number of parameters and size of training data are not disclosed. In the original Google paper on Gemini, it was reported to achieve similar performance to GPT 3.5 Turbo.
@@
in contrast, is an open-source mixture-of-experts model, consisting of eight 7B parameter models (Mistral AI team, 2023). It has been reported to achieve comparable accuracy to GPT 3.5 Turbo on several tasks, including some examined in this paper.
@@ 2.2 Model Querying Details
All models were queried through the unified interface provided by LiteLLM [^4] https://litellm.ai/ between December 11-15, 2023. Gemini was queried through Google Vertex AI, OpenAI models through the OpenAI API, and Mixtral through the API provided by Together. [^5] https://cloud.google.com/vertex-ai/docs https://openai.com/api https://docs.together.ai/docs For reference, we also list the current pricing of each model through these APIs for 1M tokens in Table 2, which provides an approximate measure of how efficiently the models can be run.
It is also notable that in some cases Gemini Pro blocks some questions, particularly in the case of potentially illegal or sensitive material. Responses were blocked for some portion of the testing examples, so we treated these examples as incorrect. For each task where a significant number of responses were blocked, we will discuss this effect in the corresponding section.
@@ 2.3 Evaluation Procedure
To perform a fair comparison between the models, we re-ran experiments with all models using exactly the same prompts and evaluation protocol for all evaluated models. We make this decision to ensure that all models are compared on exactly the same footing, in contrast to previous papers where these settings may differ. In general, we tried to follow both prompts and evaluators from standard repositories, either those officially released by the datasets themselves, or from the Eleuther evaluation harness (Gao et al., 2023a). These prompts generally consist of a query, input, and few-shot examples, sometimes including chain-of-thought reasoning (Wei et al., 2022). In some cases, we found it necessary to make small changes from standard practice to stably evaluate all models under consideration; all such deviations are noted below and implemented in the companion code repository.
@@ 3.2 Results and Analysis
Figure 2: Ratio of multiple-choice answers being prdicted by models
@@
First, from the overall results shown in Figure 2, we can see that Gemini Pro achieves an accuracy lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo. We saw little difference in performance using chain-of-thought prompting, likely because MMLU is mostly a knowledge-based question answering task that may not benefit significantly from stronger reasoning-oriented prompts.777 Note that our evaluation numbers for GPT 4 Turbo (80.5%) are slightly worse than those from GPT 4 reported by OpenAI (2023) (86.4%). This drop could likely be attributed to the “Turbo” model, but is also possibly the result of slight differences in prompting or generation methods.
Based on this overall result, we next dive a bit deeper. One first notable point is that all questions in MMLU are multiple-choice with 4 potential answers ordered A through D. In Figure 2, we show the ratio of the number of times each model selects each multiple choice answer. From this figure, we can see that Gemini has a very skewed label distribution, biased towards selecting the final choice of “D”, which contrasts to the result of the GPT model, which is more balanced. This may indicate that Gemini has not been heavily instruction-tuned towards solving multiple-choice questions, which can cause models to be biased with respect to answer ordering (Tjuatja et al., 2023).
Next, we examine each subtask’s performance. Figure 3 illustrates each model’s performance on selected representative tasks. We notice that Gemini Pro underperforms on most tasks compared to GPT 3.5. Chain-of-thought prompting decreases the variance across the subtasks.
Further, we dig deeper into the tasks where Gemini Pro underperforms/outperforms GPT 3.5 the most. From Figure 4, we can observe that Gemini Pro falls behind GPT 3.5 on human_sexuality (social sciences), formal_logic (Humanities), elementary_mathematics (STEM), and professional_medicine (specialized domains). For the two tasks where Gemini Pro outperformed GPT 3.5 Turbo, gains were marginal.
The underperformance of Gemini Pro on particular tasks can be attributed to two reasons. First, as previously mentioned in subsection 2.2, in some cases Gemini fails to return an answer. In most MMLU sub-tasks, the API response rate was greater than 95%, but two had notably low response rates: moral_scenarios at 85% and human_sexuality at 28%. This indicates that low performance on some tasks can be attributed to content filters on the input. Second, Gemini Pro performed somewhat more poorly at the basic mathematical reasoning necessary to solve the formal_logic and elementary_mathematics tasks, which we examine further in Section 4.
@@ 4.2 Results and Analysis
For the reasoning tasks, we report the overall performance, performance by question complexity, and performance by BIG-Bench sub-task.
Figure 6: Overall accuracy on BIG-Bench-Hard
@@
Figure 6: Overall accuracy on BIG-Bench-Hard
Figure 7: Accuracy by question length on BIG-Bench-Hard
First, we illustrate the overall accuracy in Figure 7, we can see that Gemini Pro achieves an accuracy slightly lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo. In contrast, the Mixtral model achieves much lower accuracy.
Based on this overall result, let us dig a little bit deeper into why Gemini might be underperforming. First, we examined accuracy by the length of the question, as detailed in Figure 7. We found that Gemini Pro underperformed on longer, more complex questions while the GPT models were more robust to this. This was particularly the case for GPT 4 Turbo, which showed very little degradation even on longer questions, indicating an impressively robust ability to understand longer and more complex queries. GPT 3.5 Turbo fell in the middle with respect to this robustness. Mixtral was notably stable with respect to question length, but had low accuracy overall.
Next we look at whether there are variations in accuracy by the specific task in BIG-Bench-Hard. Below, we list the tasks where GPT 3.5 Turbo outperformed Gemini Pro by the largest amount.
Figure 8: Tasks where GPT 3.5 Turbo excels over Gemini Pro
We can notice that Gemini Pro is particularly bad at the ‘tracking_shuffled_objects’ tasks. These tasks involve keeping track of who has certain objects as they are traded among people, and Gemini Pro often has difficulty keeping the order straight (as Figure 8).
@@
In some tasks, namely multistep_arithmetic_two, salient_translation_error_detection, snarks, disambiguition_qa and two of tracking_shuffled_objects tasks, Gemini Pro even performs worse than the Mixtral model.
However, there were a few tasks where Gemini Pro outperformed GPT 3.5 Turbo. The Figure 9 shows the six tasks where Gemini Pro outperformed GPT 3.5 Turbo by the largest amount. These were heterogeneous and included those that required world knowledge (sports_understanding), manipulating stacks of symbols (dyck_languages), sorting words in alphabetical order (word_sorting), and parsing tables (penguins_in_a_table), among others.
Figure 9: Tasks where Gemini Pro excels over GPT 3.5 Turbo
@@
We further investigate the robustness of LLMs across different answer types in the Figure below. We can see that Gemini Pro shows the worst performance in Valid/Invalid answer type which falls under the task formal_fallacies. Interestingly 68.4% of questions from this task were blocked by. However, Gemini outperformed all GPT models as well as Mixtral by a significant margin on Other answer types (consisting of the word_sorting and dyck_language tasks) which follows a similar line of findings as above i.e., Gemini is particularly good at word rearrangement and producing symbols in the correct order. Also for MCQ answers, 4.39% questions were blocked Gemini Pro, and while GPT models excel in this genre, Gemini struggles to compete with them.
In sum, there did not seem to be a particularly strong trend in which tasks one model performed better than the other, so when performing general-purpose reasoning tasks it may be worth trying both the Gemini and GPT models before making a decision on which to use.
@@ 5.1 Experimental Details
We consider standard 8-shot chain-of-thought prompts (Gao et al., 2023a; Wei et al., 2022) where each question in few-shot prompting is associated with a chain of thought for generating the corresponding answer. We evaluate all LLMs via greedy decoding using a temperature of 0.
@@
In evaluation, we make a slight modification to the standard evaluation protocol in the Eleuther harness, which consisted of matching the words “The answer is” followed by a numerical output. We found that all evaluated models had a tendency to output the correct answer even when this specific phrase was not present. To mitigate this, we simply taking the last number of the generated text as the answer of the question, which resulted in higher accuracy overall.
@@ 5.2 Results and Analysis
First, looking at overall results in the Figure 11, we can see that Gemini Pro achieves an accuracy slightly lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo on the GSM8K, SVAMP and ASDIV tasks, which all contain diverse language patterns. For the MAWPS task, all models achieve more than 90% accuracy, although Gemini Pro is still slightly worse than GPT models. Interestingly in this task GPT 3.5 Turbo outperforms GPT 4 Turbo by a close margin. In contrast, the Mixtral model achieves much lower accuracy compared to others.
@@
Figure 13: GSM8K accuracy by chain-of-thought length
Similarly to Section 4 we break down the results to observe the robustness of each model to question length in Figure 12. As with the reasoning tasks on BIG-Bench Hard, we see a drop-off on longer questions. As before, GPT 3.5 Turbo outperforms Gemini Pro on shorter questions, but drops off more quickly, with Gemini Pro achieving similar (but still slightly inferior) accuracy on longer questions.
@@
Additionally, we observe the accuracy of the LLMs when the answer requires longer chains of thought. As shown in Figure 13, GPT 4 Turbo is very robust even when using long reasoning chains, where GPT 3.5 Turbo, Gemini Pro and Mixtral struggle with increasing COT lengths. In this analysis we also find that Gemini Pro is superior to GPT 3.5 Turbo in the most complex examples where the COT length is over 100, but underperforms in the shorter examples.
Finally, we investigate the accuracy of the compared models in generating answers with varying numbers of digits. We create three buckets based on the number of digits in the answer, 1, 2, or 3+ (except for the MAWPS task which does not have answers more than two digits). As shown in Figure 14, GPT 3.5 Turbo appears to be more robust to multi-digit math problems, where Gemini Pro degrades somewhat more on problems with more digits.

### 24 Dec 2023 v2

@@ Abstract
The recently released Google Gemini class of models are the first to comprehensively report results that rival the OpenAI GPT series across a wide variety of tasks. In this paper, we do an in-depth exploration of Gemini’s language abilities, making two contributions. First, we provide a third-party, objective comparison of the abilities of the OpenAI GPT and Google Gemini models with reproducible code and fully transparent results. Second, we take a closer look at the results, identifying areas where one of the two model classes excels. We perform this analysis over 10 datasets testing a variety of language abilities, including reasoning, answering knowledge-based questions, solving math problems, translating between languages, generating code, and acting as instruction-following agents. From this analysis, we find that Gemini Pro achieves accuracy that is close but slightly inferior to the corresponding GPT 3.5 Turbo on all English-language tasks that we benchmarked, but find that Gemini Pro excels in translation into other languages for the languages that it supports. We further provide explanations for some of the under-performing tasks, including failures in mathematical reasoning with many digits, sensitivity to multiple-choice answer ordering, and others. We also identify areas where Gemini Pro demonstrates comparably high performance, such as handling longer and more complex reasoning chains. Code and data for reproduction can be found at https://github.com/neulab/gemini-benchmark
@@ 1 Introduction
Gemini is the most recent in a series of large language models released by Google DeepMind (Gemini Team, 2023). It is notable in particular because the results reported by the Gemini team are the first to rival the OpenAI GPT model series (Brown et al., 2020) across a wide variety of tasks. Specifically, Gemini’s “Ultra” version reportedly outperforms GPT-4 on a wide variety of tasks, while Gemini’s “Pro” version is reportedly comparable to GPT-3.5 (Gemini Team, 2023). Despite the potential impact of these results, the exact evaluation details and model predictions have not been released, limiting the ability to reproduce, inspect, and analyze the results and their implications in detail.
@@
A summary of our main results can be found in Table 1. In sum, we found that across all tasks, as of this writing (December 24, 2023), Gemini’s Pro model achieved comparable but slightly inferior accuracy compared to the current version of OpenAI’s GPT 3.5 Turbo for all English tasks, but superior ability to translate into other languages that it supports. Mixtral was competitive with the Gemini and GPT models for Knowledge-based QA and Mathematics tasks, but fell short in more complex tasks. In the following sections, we will detail our experimental methodology (Section 2) and then perform an in-depth description and analysis of the results on each task. Each analysis is accompanied by an online results browser using Zeno (Cabrera et al., 2023), [^2] https://zenoml.com which can be accessed through the [Uncaptioned image] images in this PDF. All results and code for reproduction can be found at https://github.com/neulab/gemini-benchmark.
@@
Knowledge-based QA MMLU (5-shot) 65.22 67.75 80.48 68.81
MMLU (CoT) 62.09 70.07 78.95 59.57
Reasoning BIG-Bench-Hard 67.53 71.02 83.90 60.76
Mathematics GSM8K 76.42 78.01 92.72 71.65
SVAMP 81.10 82.30 92.60 81.60
ASDIV 85.31 89.07 92.75 83.16
MAWPS 96.50 98.00 98.67 96.00
Code Generation HumanEval 59.76 74.39 76.83 45.12
ODEX 39.86 52.62 45.79 40.55
Machine Translation FLORES (5-shot) Unblocked 53.31 52.43 54.00 40.97
FLORES (5-shot) All 21.68 40.00 48.24 30.27
Web Agents WebArena 7.12 8.87 14.90 1.39
Table 1: Main results of our benchmarking. The best model is listed in bold, and the second best model is underlined.
@@ 2.1 Models Tested
is the second largest model in the Gemini Series, next to the largest Gemini Ultra. [^3] Gemini Ultra is not yet publicly available, and thus we do not test it in the current version of this paper. The model is based on the Transformer (Vaswani et al., 2017) architecture and was trained multimodally over videos, text, and images. The number of parameters and size of training data are not disclosed. In the original Google paper on Gemini, it was reported to achieve similar performance to GPT 3.5 Turbo.
@@
in contrast, is an open-source mixture-of-experts model, where each feedforward block picks from a set of 8 distinct groups of parameters and uses two to process the token (Mistral AI team, 2023). It has been reported to achieve comparable accuracy to GPT 3.5 Turbo on several tasks, including some examined in this paper. We use the mistralai/Mixtral-8x7b-Instruct-v0.1 version of the model.
@@ 2.2 Model Querying Details
All models were queried through the unified interface provided by LiteLLM [^4] https://litellm.ai/ between December 11-22, 2023. Gemini was queried through Google Vertex AI, OpenAI models through the OpenAI API, and Mixtral through the API provided by Together. [^5] https://cloud.google.com/vertex-ai/docs https://openai.com/api https://docs.together.ai/docs For reference, we also list the current pricing of each model through these APIs for 1M tokens in Table 2, which provides an approximate measure of how efficiently the models can be run.
It is also notable that in some cases Gemini Pro, by default, has safety features [^6] https://cloud.google.com/vertex-ai/docs/generative-ai/multimodal/configure-safety-attributes that block some questions, particularly in the case of potentially illegal or sensitive material. For analysis in this paper, we disabled these safety settings, but in some cases discuss the effect on measured accuracy (contrasting with gemini-pro-filtered, a model with these safety filters enabled).
@@ 2.3 Evaluation Procedure
To perform a fair comparison between the models, we re-ran experiments with all models using exactly the same prompts and evaluation protocol for all evaluated models. We make this decision to ensure that all models are compared on exactly the same footing, in contrast to previous papers where these settings may differ. In general, we tried to follow both prompts and evaluators from standard repositories, either those officially released by the datasets themselves, or from the Eleuther evaluation harness (Gao et al., 2023a). We also personally communicated with the Gemini team, and in some cases followed their recommended prompts for evaluating the models, in the cases where these prompts provided uniformly superior to the standard prompts over all evaluated model classes. These prompts generally consist of a query, input, and few-shot examples, sometimes including chain-of-thought reasoning (Wei et al., 2022). In some cases, we found it necessary to make small changes from standard practice to stably evaluate all models under consideration; all such deviations are noted below and implemented in the companion code repository.
@@ 3.2 Results and Analysis
Figure 2: Ratio of multiple-choice answers being predicted by models
@@
First, from the overall results shown in Figure 2, we can see that Gemini Pro achieves an accuracy close to, but slightly lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo. MMLU is the strongest task for the Mixtral model, with it besting both Gemini Pro and GPT 3.5 Turbo. We saw some degradation of performance using chain-of-thought prompting, likely because MMLU is mostly a knowledge-based question answering task that may not benefit significantly from reasoning-oriented prompts.888 Note that our evaluation numbers are slightly lower than those reported in the Gemini (Gemini Team, 2023) and Mixtral (Mistral AI team, 2023) technical reports. We attribute this to sensitivity to prompts – other prompts may achieve somewhat higher accuracy overall.
Based on this overall result, we next dive a bit deeper. One first notable point is that all questions in MMLU are multiple-choice with 4 potential answers ordered A through D. In Figure 2, we show the ratio of the number of times each model selects each multiple choice answer. From this figure, we can see that Gemini has a very skewed label distribution, biased towards selecting the final choice of “D”, which contrasts to the result of the other models, which are more balanced. This may indicate that Gemini has not been heavily instruction-tuned towards solving multiple-choice questions, which can cause models to be biased with respect to answer ordering (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023; Tjuatja et al., 2023).
Next, we examine each subtask’s performance. Figure 3 illustrates each model’s performance on selected representative tasks. We notice that Gemini Pro underperforms on most tasks compared to GPT 3.5 Turbo. Chain-of-thought prompting decreases the variance across the subtasks.
Further, we dig deeper into the tasks where Gemini Pro underperforms/outperforms GPT 3.5 Turbo the most. From Figure 4, we can observe that Gemini Pro falls behind GPT 3.5 Turbo on human_sexuality, marketing abstract_algebra, and miscellaneous. In contrast, Gemini Pro excels at both college_biology and high_school_biology, as well as high_school_macroeconomics and security_studies.
One important thing to note is that, as previously mentioned in subsection 2.2, Gemini Pro’s safety features can have a significant effect on overall performance. All results reported above are Gemini Pro with safety filtering disabled, but when the features are enabled the response rate and corresponding performance can drop. In most MMLU sub-tasks, the API response rate was greater than 95%, but two had notably low response rates: moral_scenarios at 85% and human_sexuality at 28%.
@@ 4.2 Results and Analysis
For the reasoning tasks, we report the overall performance, performance by question complexity, performance by answer types and performance by BIG-Bench sub-task.
Figure 6: Overall accuracy on BIG-Bench Hard
@@
Figure 6: Overall accuracy on BIG-Bench Hard
Figure 7: Accuracy by question length on BIG-Bench Hard
First, we illustrate the overall accuracy in Figure 7, we can see that Gemini Pro achieves an accuracy slightly lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo. In contrast, the Mixtral model achieves somewhat lower accuracy than Gemini Pro.
Based on this overall result, let us dig a little bit deeper into why Gemini might be underperforming. First, we examined accuracy by the length of the question, as detailed in Figure 7. We found that Gemini Pro underperformed on longer, more complex questions while the GPT models were more robust to this. This was particularly the case for GPT 4 Turbo, which showed very little degradation even on longer questions, indicating an impressively robust ability to understand longer and more complex queries. GPT 3.5 Turbo fell in the middle with respect to this robustness. Mixtral had low accuracy overall compared to the Gemini GPT models but was notably stable with respect to question length until length 70.
Next we look at whether there are variations in accuracy by the specific task in BIG-Bench Hard. In 7(a), we list the tasks where GPT 3.5 Turbo outperformed Gemini Pro by the largest amount.
(a) Tasks where GPT 3.5 excels over Gemini Pro
(b) Tasks where Gemini Pro excels over GPT 3.5
Figure 8: Tasks where Gemini Pro and GPT 3.5 prevail on BBH
We can notice that Gemini Pro is particularly bad at the ‘tracking_shuffled_objects’ tasks. These tasks involve keeping track of who has certain objects as they are traded among people, and Gemini Pro often has difficulty keeping the order straight (as Table 3).
@@
However, there were a few tasks where Gemini Pro outperformed GPT 3.5 Turbo. The 7(b) shows the seven tasks where Gemini Pro outperformed GPT 3.5 Turbo by the largest amount. These were heterogeneous and included those that required world knowledge (sports_understanding), manipulating stacks of symbols (dyck_languages), sorting words in alphabetical order (word_sorting), detecting sarcasm (snarks) and parsing tables (penguins_in_a_table), among others.
Tasks that mostly require natural language understanding (i.e., semantic understanding, name disambiguation, entity resolution, grammar rules, or sarcasm/humor detection), namely salient_translation_error_detection, snarks, hyperbaton, disambiguition_qa, and ruin_names, the Mixtral model performed particularly well, often outperforming both Gemini and GPT 3.5 Turbo (as in Figure 9).
We further investigate the robustness of LLMs across different answer types in the Figure below. We can see that Gemini Pro shows the worst performance in Valid/Invalid answer type which falls under the task formal_fallacies representing logical deduction from a given context. However, Gemini outperformed all GPT models as well as Mixtral on Other answer types (consisting of the word_sorting and dyck_language tasks) which follows a similar line of findings as above i.e., Gemini is particularly good at word rearrangement and producing symbols in the correct order. Also for MCQ and Digit answers, GPT models excel in this genre while Gemini and Mixtral struggles to compete with them.
Figure 9: Tasks where Mixtral excels over GPT 3.5 Turbo and Gemini
@@
In sum, there did not seem to be a particularly strong trend in which tasks one model performed better than the other, so when performing general-purpose reasoning tasks it may be worth trying both the Gemini and GPT models before making a decision on which to use. On the other hand, between Gemini and Mixtral, Mixtral is more reliable in multi-variable reasoning and natural language understanding tasks, among other.
@@ 5.1 Experimental Details
We consider standard 8-shot chain-of-thought prompts (Gao et al., 2023a; Wei et al., 2022) where each question in few-shot prompting is associated with a chain of thought for generating the corresponding answer. All models use greedy decoding using a temperature of 0.
@@
In evaluation, we make a slight modification to the standard evaluation protocol in the Eleuther harness, which consisted of matching the words “The answer is” followed by a numerical output. We found that all evaluated models had a tendency to output the correct answer even when this specific phrase was not present. To mitigate this, we are simply taking the last number of the generated text as the answer to the question, which resulted in higher accuracy overall.
@@ 5.2 Results and Analysis
First, looking at overall results in the Figure 11, we can see that Gemini Pro achieves an accuracy slightly lower than that of GPT 3.5 Turbo, and much lower than that of GPT 4 Turbo on the GSM8K, SVAMP and ASDIV tasks, which all contain diverse language patterns. For the MAWPS task, all models achieve more than 90% accuracy, although Gemini Pro is still slightly worse than the other models. In contrast, the Mixtral model achieves slightly lower accuracy compared to Gemini Pro except on the SVAMP task.
Similarly to Section 4 we break down the results to observe the robustness of each model to question length in Figure 12. As with the reasoning tasks on BIG-Bench Hard, we see a drop-off on longer questions. As before, GPT 3.5 Turbo outperforms Gemini Pro on shorter questions, but drops off more quickly, with Gemini Pro achieving similar (but still slightly inferior) accuracy on longer questions - except the MAWPS task where Gemini excels over GPT 3.5 on longer questions. Mixtral’s degradation is slightly worse than Gemini’s except for SVAMP.
@@
Additionally, we observe the accuracy of the models when the answer requires longer chains of thought. As shown in Figure 14, GPT 4 Turbo is very robust even when using long reasoning chains, where GPT 3.5 Turbo, Gemini Pro, and Mixtral struggle with increasing COT lengths. In this analysis, we also find that Gemini Pro is superior to GPT 3.5 Turbo in the most complex examples where the COT length is over 100, but underperforms in the shorter examples. In contrast, Mixtral is more affected by longer chain-of-thoughts compared to other models, showing the lowest performance in the most complex examples.
Figure 14: GSM8K accuracy by chain-of-thought length
Finally, we investigate the accuracy of the compared models in generating answers with varying numbers of digits. We create three buckets based on the number of digits in the answer, 1, 2, or 3+ (except for the MAWPS task which does not have answers of more than two digits). As shown in Figure 13, GPT 3.5 Turbo appears to be more robust to multi-digit math problems, where Gemini Pro and Mixtral degrades somewhat more on problems with more digits.
To summarize, GPT 4 Turbo is the best model across all math word problem tasks, showing more than 90% accuracy in all tasks. In contrast, Gemini Pro and Mixtral lag behind all GPT models in this domain, while Gemini shows slightly better reasoning ability than Mixtral.